### v1

# OceanSplat: Object-aware Gaussian Splatting with Trinocular View Consistency for Underwater Scene Reconstruction

Abstract
We introduce OceanSplat, a novel 3D Gaussian Splatting-based approach for accurately representing 3D geometry in underwater scenes. To overcome multi-view inconsistencies caused by underwater optical degradation, our method enforces trinocular view consistency by rendering horizontally and vertically translated camera views relative to each input view and aligning them via inverse warping. Furthermore, these translated camera views are used to derive a synthetic epipolar depth prior through triangulation, which serves as a self-supervised depth regularizer. These geometric constraints facilitate the spatial optimization of 3D Gaussians and preserve scene structure in underwater environments. We also propose a depth-aware alpha adjustment that modulates the opacity of 3D Gaussians during early training based on their $z$-component and viewing direction, deterring the formation of medium-induced primitives. With our contributions, 3D Gaussians are disentangled from the scattering medium, enabling robust representation of object geometry and significantly reducing floating artifacts in reconstructed underwater scenes. Experiments on real-world underwater and simulated scenes demonstrate that OceanSplat substantially outperforms existing methods for both scene reconstruction and restoration in scattering media.
Project Pagehttps://oceansplat.github.io
Introduction
Underwater scene reconstruction is essential for several marine robotics tasks, including seafloor mapping (kapoutsis2016real; liu2020underwater; joshi2022underwater), ecological monitoring (marre2019monitoring; girdhar2023curee), and subsea infrastructure inspection (fang2023integration; huang2025visual). While visual data can generally support robotic operations effectively, the underwater optical properties, such as wavelength-dependent attenuation (akkaynak2017space), scattering (mcglamery1980computer; jaffe2002computer; chen2021underwater), and low illumination (marques2020l2uwe), significantly degrade perceptual cues (yu2023task; zhang2024dcgf) and hinder the deployment of vision-based autonomous systems (de2021impact). Consequently, achieving geometric understanding and spatial representation from this low-quality imagery is essential for practical operation of autonomous platforms. Although advances in remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) have greatly enabled the acquisition of large-scale underwater imagery (wynn2014autonomous), leveraging this data for scene representation remains highly challenging due to adverse visual conditions (beall20103d; huang2025visual).
Figure 1: OceanSplat overcomes scattering and attenuation effects through trinocular view consistency, preserving object geometry and enabling high-quality underwater 3D reconstruction.
To address these issues in underwater perception, the use of Neural Radiance Fields (NeRF) (mildenhall2020nerf) and 3D Gaussian Splatting (3DGS) (kerbl20233d), which have shown strong capability in 3D reconstruction, has emerged as a promising direction. However, these methods rely on assumptions tailored to in-air image acquisition, motivating recent research to integrate underwater image formation model (akkaynak2019sea).
NeRF-based methods (sethuraman2023waternerf; zhang2023beyond; levy2023seathru; ramazzina2023scatternerf; tang2024neural) extend their volumetric rendering framework by integrating light-transmitting physics and scattering phenomena. However, these methods implicitly represent both object and backscatter components, which impedes accurate geometric representation and also results in slow rendering speed. More recently, 3DGS-based methods (yang2024seasplat; li2024watersplatting; wang2024uw) have been adapted for underwater environments, benefiting from their superior rendering performance and fast training. Despite notable progress in these studies, medium intensity is often absorbed into the 3D Gaussians, leading to numerous floating artifacts and degrading reconstruction quality.
Unlike previous approaches, our goal is to robustly recover 3D geometry in underwater scenes subject to severe scattering and attenuation, without depending on external geometric supervision. To this end, we propose OceanSplat: Object-aware Gaussian Splatting for geometrically consistent underwater scene reconstruction. Our method enforces trinocular view consistency across translated camera views along two orthogonal axes, resolving multi-view inconsistencies and effectively guiding 3D Gaussians to align with the scene structure. Additionally, we derive a synthetic epipolar depth prior via triangulation across the translated views, and leverage it for self-supervised depth regularization. Furthermore, we apply a depth residual loss that constrains the $z$-component of each 3D Gaussian toward the alpha-blended depth to reduce floating artifacts. We also propose a depth-aware alpha adjustment that regulates the transparency of 3D Gaussians during the early stages of training based on their $z$-component and viewing direction, thereby discouraging the formation of medium-induced primitives. As shown in Figure 1, our method improves geometric accuracy for object representation under challenging scattering conditions by preventing the entanglement of 3D Gaussians with the scattering medium. Our key contributions are outlined in the following points:
  • We propose OceanSplat, a novel underwater scene reconstruction method leveraging 3D Gaussian Splatting that employs trinocular view consistency to effectively preserve scene geometry in scattering media.
  • We introduce a synthetic epipolar depth prior derived through triangulation from translated camera views, enabling self-supervised geometric regularization.
  • We propose a depth-aware alpha adjustment that regulates the opacity of 3D Gaussians during the early training stage, suppressing medium-induced primitives.
  • Our approach surpasses contemporary state-of-the-art methods in both qualitative and quantitative evaluations across real-world and simulated datasets.
Figure 2: Overview of OceanSplat. We enforce trinocular view consistency by inverse warping rendered images from two translated camera poses to guide the spatial optimization of 3D Gaussians. From these poses, we derive a synthetic epipolar depth prior via triangulation, which provides self-supervised geometric constraints. Additionally, depth-aware alpha adjustment suppresses erroneous 3D Gaussians early and aligns rendered depth with the Gaussian $z$-component to prevent floating artifacts.
Related Work
Underwater Scene Reconstruction
Underwater scene reconstruction aims to represent precise scene structure and restore true color despite severe light attenuation and scattering. Initial approaches jointly estimated medium-dependent attenuation during reconstruction (skinner2017automatic) and exploited global context with active labeling for sonar-based modeling (debortoli2018real). Subsequent works (sethuraman2023waternerf; zhang2023beyond; tang2024neural) embed physics-based underwater light transport and scattering into NeRF (mildenhall2020nerf) to jointly recover geometry and true color appearance. Seathru-NeRF (levy2023seathru) leverages the SeaThru image formation model (akkaynak2019sea) within NeRF to successfully achieve novel view synthesis in real-world underwater environments. ScatterNeRF (ramazzina2023scatternerf) uses inverse rendering, while DecoNeRF (zhang2025decoupling) employs pseudo-labeling to disentangle scattering media from scene content and restore clear views in foggy scenes. Although these methods significantly improve reconstruction quality, they are often hindered by slow rendering speed and high memory consumption. Recently, 3DGS (kerbl20233d) has gained increasing attention for underwater scene reconstruction. SeaSplat (yang2024seasplat) incorporates underwater physics into 3D Gaussians to improve color restoration and structural fidelity, while WaterSplatting (li2024watersplatting) combines implicit medium modeling with explicit object representation for efficient, high-quality reconstruction. UW-GS (wang2024uw) further enhances geometry using pseudo-depth priors from foundation models (yang2024depth). Despite these advances, existing 3DGS-based methods struggle to represent geometry-consistent underwater scenes.
In contrast, our method enforces a trinocular view consistency constraint that enhances the spatial coherence of 3D Gaussians, and leverages self-supervised depth regularization based on epipolar geometry to achieve geometrically accurate underwater scene reconstruction.
Multi-View Stereo
Multi-view stereo (MVS) algorithms reconstruct detailed 3D geometry by leveraging images acquired from multiple calibrated viewpoints. Early approaches employed voxelization (kutulakos2000theory), dense point clouds (furukawa2009accurate), and depth maps (galliani2015massively; schonberger2016structure). In the deep learning era, MVSNet (yao2018mvsnet) introduced cost volumes to form geometry-aware 3D representations, greatly enhancing accuracy. Building upon these advancements, MVS has been integrated into neural rendering (mildenhall2020nerf; kerbl20233d), where NeRF-based methods (chen2021mvsnerf; wang2021neus; fu2022geo; wu2025sparis) utilize these concepts for surface reconstruction, while 3DGS-based approaches (chen2023neusg; guedon2024sugar; huang20242d; chung2024depth; chen2024pgsr; chen2024mvsplat; liu2024mvsgaussian; li2024dngaussian; huang2025fatesgs) harness multi-view relationships to enhance geometric fidelity. Among these, recent efforts (safadoust2024self; han2024binocular) have demonstrated promising results by constructing virtual stereo pairs to enforce binocular stereo consistency, effectively regularizing 3D Gaussians.
To better address geometric ambiguities inherent to scattering media, we extend binocular stereo vision to a trinocular configuration by introducing a vertically translated camera, facilitating spatial optimization of 3D Gaussians.
Preliminaries
3D Gaussian Splatting
3D Gaussian Splatting (kerbl20233d) models a scene with anisotropic 3D Gaussian primitives, each parameterized by a center position $\mu\in\mathbb{R}^{3}$, covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, color $c$, opacity $\alpha$, and view-dependent appearance modeled by spherical harmonics. Each 3D Gaussian defines its spatial influence at a point $\mathbf{X}\in\mathbb{R}^{3}$ centered at $\mu$ as:
$G(\mathbf{X})=e^{-\tfrac{1}{2}(\mathbf{X}-\mu)^{\top}\Sigma^{-1}(\mathbf{X}-\mu)}.$ (1)
For rendering, 3D Gaussians are projected onto 2D screen space and composited via alpha-blending of $N$ primitives that are sorted by depth and overlap at each pixel:
$C=\sum_{i=1}^{N}T_{i}\cdot\alpha_{i}\cdot c_{i},\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
where $C$, NN, and $T_{i}$ denote the rendered color, the number of 3D Gaussians intersected by the ray, and the transmittance, respectively. Then, 3D Gaussians are optimized via an L1 loss combined with D-SSIM on the rendered images.
Underwater Image Formation Model
To render a 3D underwater scene, the image formation is guided by the revised underwater image formation model (akkaynak2018revised), which decomposes the observed image into a direct component attenuated by the medium with respect to depth $z$, and a backscatter component accumulated along the viewing direction, as follows:
$C=C^{\text{obj}}\cdot e^{-\sigma^{\text{attn}}\cdot z}+C^{\infty}\cdot\left(1-e^{-\sigma^{\text{bs}}\cdot z}\right),$ (3)
where $^{\text{obj}}, CC^{\infty}$, $\sigma^{\text{attn}}$, and $\sigma^{\text{bs}}$ denote the unattenuated object color, the asymptotic backscatter color, the attenuation coefficient, and the backscattering coefficient, respectively.
Methodology
Figure 2 shows the overall architecture of our method. We begin by initializing 3D Gaussians from a collection of RGB images $I\in\mathbb{R}^{H\times W\times 3}$, along with their corresponding camera intrinsics $K\in\mathbb{R}^{3\times 3}$ and extrinsic parameters $P\in\mathbb{R}^{4\times 4}$, which represent the homogeneous transformation matrix, recovered using Structure-from-Motion (SfM) (schonberger2016structure). We model medium properties along the ray $\vec{\mathbf{r}}$ using an MLP $\phi_{\text{med}}$, following (li2024watersplatting), which directly predicts attenuation $\sigma^{\text{attn}}$, backscattering $\sigma^{\text{bs}}$, and medium color $c^{\text{med}}$ as follows:
$\sigma^{\text{attn}},\ \sigma^{\text{bs}},\ c^{\text{med}}=\phi_{\text{med}}(\vec{\mathbf{r}}).$ (4)
Subsequently, by integrating Eq. (2) with Eq. (3), the contributed colors of the objects and medium along the ray in underwater scenes are formulated as follows:
$C^{\text{obj}}=\sum_{i=1}^{N}T^{\text{obj}}_{i}\cdot\alpha_{i}\,\cdot c_{i}\,\cdot e^{-\sigma^{\text{attn}}z_{i}},T^{\text{obj}}_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (5)
$C^{\infty}=T^{\text{obj}}_{N}\cdot c_{\text{med}}\cdot e^{-\sigma^{\text{bs}}z_{N}},$ (6)
$C^{\text{med}}=\sum_{i=1}^{N}\left[T^{\text{obj}}_{i}\cdot c_{\text{med}}\cdot(e^{-\sigma^{\text{bs}}z_{i-1}}-e^{-\sigma^{\text{bs}}z_{i}})\right]+C^{\infty},$ (7)
where $C^{\text{obj}}$, $^{\text{med}}, CC^{\infty}$, and $T^{\text{obj}}$ denote the pixel-wise colors of the object, medium, background medium, and represent the transmittance of 3D Gaussians, respectively. Here, the object refers to physical structures submerged in water, while the medium represents the water itself. Based on this rendering framework, we describe our proposed methods for robust 3DGS-based underwater scene reconstruction.
Trinocular View Consistency
We propose a trinocular stereo framework for 3DGS to improve object representation in scattering media. While prior works (safadoust2024self; han2024binocular) use only binocular stereo consistency, we show that a vertically translated viewpoint provides orthogonal constraints that complement horizontal stereo for 3D Gaussians. As shown in Figure 3, view-dependent sampling in alpha-blending often causes inconsistencies under minor viewpoint changes. Furthermore, previous stereo matching studies (okutomi1993multiple; imran2020unsupervised; shamsafar2021tristereonet) have shown that multi-baseline stereo surpasses single-baseline methods in accuracy and robustness. Inspired by these observations, we enforce trinocular view consistency with orthogonal baselines of varying lengths to better regularize the spatial position of 3D Gaussians.
Figure 3: Pixel-wise color contributions across translated camera views in 3D Gaussian Splatting.
From a given camera pose $P_{c}$, we obtain the rasterized object and medium components, denoted by $I^{\text{obj}}_{c}$ and $I^{\text{med}}_{c}$, respectively. Two additional camera poses are defined by translating $P_{c}$ along the horizontal $h$ and vertical $v$ axes as:
$P_{h}=\begin{bmatrix}\mathbb{I}&\textbf{t}_{h}\\ \textbf{0}^{\top}&1\end{bmatrix}P_{c},\quad P_{v}=\begin{bmatrix}\mathbb{I}&\textbf{t}_{v}\\ \textbf{0}^{\top}&1\end{bmatrix}P_{c},$ (8)
where $\mathbb{I}$ is the $3\times 3$ identity matrix, $\mathbf{t}_{h}=(b_{h},0,0)^{\top}$ and $\mathbf{t}_{v}=(0,b_{v},0)^{\top}$ are translation vectors defining the horizontally and vertically translated camera poses $P_{h}$ and $P_{v}$, respectively, with $b_{h},b_{v}\in\mathbb{R}$ denoting the baseline distances. Using these poses, we render two images $I^{\text{obj}}_{h}$ and $I^{\text{obj}}_{v}$ from the poses $P_{h}$ and $P_{v}$. We then apply disparity-based inverse warping to the rendered images. Let $D_{c}\in\mathbb{R}^{H\times W\times 1}$ be the rendered depth map from $P_{c}$. Based on stereo geometry, the horizontal and vertical disparity maps are computed as:
$d_{h}(x,y)=\frac{f_{h}\cdot b_{h}}{D_{c}(x,y)},\quad d_{v}(x,y)=\frac{f_{v}\cdot b_{v}}{D_{c}(x,y)},$ (9)
where $f_{h},f_{v}\in\mathbb{R}$ denote the focal lengths in the horizontal and vertical directions. The resulting disparity maps $d_{h},d_{v}\in\mathbb{R}^{H\times W\times 1}$ represent the pixel-wise displacements needed to align the translated views with the central view, and $(x,y)$ denotes the spatial pixel coordinates. Then, the rendered images from the translated views, $I^{\text{obj}}_{h}$ and $I^{\text{obj}}_{v}$, are inverse-warped using disparity maps $d_{h}$ and $d_{v}$, respectively, in the direction opposite to the virtual camera shift, aligning them with the central image rendered from $P_{c}$. The resulting inverse warped images $I^{\text{obj}}_{h\rightarrow c},I^{\text{obj}}_{v\rightarrow c}\in\mathbb{R}^{H\times W\times 3}$ are defined as:
$\left\{\begin{array}[]{l}I^{\text{obj}}_{h\rightarrow c}(x,y)=I^{\text{obj}}_{h}(x-d_{h}(x,y),\ y),\\ I^{\text{obj}}_{v\rightarrow c}(x,y)=I^{\text{obj}}_{v}(x,\ y-d_{v}(x,y)),\end{array}\right$. (10)
where $d_{h},d_{v}\in\mathbb{R}^{H\times W}$ denote the horizontal and vertical disparities. To reconstruct the complete scene, the shifted object images are composited with the medium component $I^{\text{med}}_{c}$ as:
$\left\{\begin{array}[]{l}I_{c}=I^{\text{obj}}_{c}+I^{\text{med}}_{c},\\ I_{h\rightarrow c}=I^{\text{obj}}_{h\rightarrow c}+I^{\text{med}}_{c},\\ I_{v\rightarrow c}=I^{\text{obj}}_{v\rightarrow c}+I^{\text{med}}_{c}.\end{array}\right$. (11)
Afterward, to emphasize dark regions as perceived in human vision, we follow prior studies (mildenhall2022nerf; levy2023seathru; li2024watersplatting), and employ a regularized L1 loss $L_{\text{R-L1}}$ defined as:
$L_{\text{R-L1}}(I_{1},I_{2})=\frac{1}{HW}\sum_{x,y}\left|\frac{I_{1}(x,y)-I_{2}(x,y)}{\mathsf{sg}(I_{c}(x,y))+\epsilon}\right|,$ (12)
where $\mathsf{sg}(\cdot)$ and $\epsilon$ denote the stop-gradient operator and a small constant for numerical stability, respectively. This loss enforces object consistency across rendered images and photometric alignment with the ground truth $I_{gt}$ as follows:
$L_{\text{obj-stereo}}=L_{\text{R-L1}}(I^{\text{obj}}_{h\rightarrow c},I^{\text{obj}}_{c})+L_{\text{R-L1}}(I^{\text{obj}}_{v\rightarrow c},I^{\text{obj}}_{c}),$ (13)
$L_{\text{full-stereo}}=L_{\text{R-L1}}(I_{h\rightarrow c},I_{gt})+L_{\text{R-L1}}(I_{v\rightarrow c},I_{gt}).$ (14)
In addition, we define the smoothness loss for horizontal and vertical disparities as follows:
$L_{\text{smooth}}=\frac{1}{HW}\sum_{d\in\{d_{h},d_{v}\}}\sum_{x,y}\sum_{k\in\{h,v\}}\left|\nabla_{k}d\right|\cdot e^{-\gamma\left|\nabla_{k}I_{gt}\right|},$ (15)
where $\nabla_{k}$ denotes the directional gradient along $k$-axis and $\gamma$ is a weighting factor, respectively. Finally, the overall trinocular view consistency loss is defined as:
$L_{\text{tri}}=L_{\text{obj-stereo}}+L_{\text{full-stereo}}+L_{\text{smooth}}.$ (16)
Ground Truth SeaThru-NeRF 3DGS SeaSplat WaterSplatting Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative evaluation of novel view synthesis on diverse real-world underwater 3D scenes.
Synthetic Epipolar Depth Prior
To enhance geometric alignment, we introduce a synthetic epipolar depth prior $D_{\text{epi}}$, derived via triangulation between two translated views. For reliable correspondences in triangulation, we use 3D Gaussians sampled from the intersection of the view frusta of all trinocular camera views, with opacity above $\tau_{\alpha}$. Each selected Gaussian is projected onto views $P_{h}$ and $P_{v}$ via projection matrices $\mathbf{M}_{h},\mathbf{M}_{v}\in\mathbb{R}^{3\times 4}$, producing screen-space points $\mathbf{x}_{h}^{i},\mathbf{x}_{v}^{i}\in\mathbb{R}^{2}$. Let the homogeneous coordinates be $\tilde{\mathbf{x}}_{k}^{i}=[\mathbf{x}_{k}^{i\top},1]^{\top}\in\mathbb{R}^{3}$ and $\tilde{\mathbf{X}}_{i}=[\mathbf{X}_{i}^{\top},1]^{\top}\in\mathbb{R}^{4}$. We apply epipolar geometry to the projected correspondences, which yields a homogeneous linear system $\mathbf{A}_{i}\tilde{\mathbf{X}}_{i}=\mathbf{0}$, where $\mathbf{A}_{i}\in\mathbb{R}^{4\times 4}$ stacks the constraints from both views. To solve it stably, we rewrite it in the least-squares form:
$\hat{\mathbf{X}}_{i}=\arg\min_{\mathbf{X}}\left\|\mathbf{A}_{i}^{\prime}\mathbf{X}+\mathbf{b}_{i}\right\|_{2}^{2},$ (17)
where $\mathbf{A}_{i}^{\prime}\in\mathbb{R}^{4\times 3}$ and $\mathbf{b}_{i}\in\mathbb{R}^{4}$ are obtained by separating the last column of $\mathbf{A}_{i}$. The triangulated point $\hat{\mathbf{X}}_{i}$ is transformed to the central camera view via extrinsics $\mathbf{R}_{c}\in\mathbb{R}^{3\times 3}$, $\mathbf{t}_{c}\in\mathbb{R}^{3}$, and the depth prior is defined as the $z$-component:
$D_{\text{epi}}=\left[\mathbf{R}_{c}\hat{\mathbf{X}}_{i}+\mathbf{t}_{c}\right]_{z}.$ (18)
After that, we apply an edge-aware Log-L1 loss (turkulainen2025dn) using $D_{\text{epi}}$ as follows:
$L_{\text{epi}}=\frac{1}{HW}\sum_{x,y}\sum_{k\in\{h,v\}}\log\!\left(1+\big|D_{c}^{{}^{\prime}}-D_{\text{epi}}\big|\right)\cdot e^{-\bigl|\nabla_{k}I_{c}\bigr|},$ (19)
where $D_{c}^{{}^{\prime}}$ denotes the depth sampled at the same spatial locations as $D_{\text{epi}}$. Through this approach, we enforce depth regularization in a self-supervised manner.
Method SeaThru-NeRF In-the-Wild
Curaçao Panama JapaneseGardens-Redsea IUI3-Redsea Coral Composite
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
Table 1: Quantitative evaluation of novel view synthesis performance on real-world underwater scenes compared with reproducible prior methods. 1st, 2nd, and 3rd indicate performance ranking.
Method Underwater Foggy
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Novel View Synthesis
[…]
Scene Restoration
SeaThru-NeRF-NS 12.54 0.544 0.400 14.87 0.605 0.303
[…]
Table 2: Quantitative evaluation of novel view synthesis and scene restoration performance on simulated scenes.
Depth Residual Loss
Overly dispersed 3D Gaussians along a camera ray may cause those far from the actual surface to appear as floating artifacts in novel views. To alleviate this issue, we apply a depth residual loss that aligns the rendered depth with the $z$-component of individual 3D Gaussians. This residual loss is computed as follows:
$L_{\text{res}}=\frac{1}{N^{\prime}}\sum_{i=1}^{N^{\prime}}\left|D_{c}(\mathbf{x}_{i})-z_{i}\right|,$ (20)
where $N^{\prime}$ denotes the number of 3D Gaussians located within the camera view frustum, $D_{c}(\mathbf{x}_{i})$ is the rendered depth via alpha-blending at pixel location $\mathbf{x}_{i}$, and $z_{i}$ is the camera-space $z$-component of the corresponding 3D Gaussian.
Depth-aware Alpha Adjustment
In scattering media, misplaced 3D Gaussians acquire medium-colored contributions that appear as floating artifacts in novel views. To tackle this challenge, we introduce a depth-aware alpha adjustment. During the adjustment stage $t<t_{\alpha}$, the adjusted opacity $\alpha^{\prime}_{i}$ is used for rasterization and is formulated as a weighted sum of the original opacity $\alpha_{i}$ and the depth-aware opacity $\alpha^{d}_{i}$ as follows:
$\alpha^{\prime}_{i}=(1-w)\,\alpha_{i}+w\,\alpha^{d}_{i},\qquad\alpha^{d}_{i}=\phi_{\alpha}(\alpha_{i},z_{i},\vec{\mathbf{v}}_{i}),$ (21)
where $\phi_{\alpha}$ is the MLP and $\vec{\mathbf{v}}_{i}$ is the viewing direction vector. Note that, after transition step $t_{\alpha}$, the weight $w$ is decayed to zero to eliminate inference-time overhead. This approach adaptively adjusts the transparency of each 3D Gaussian using depth and viewing direction cues, suppresses their opacity in scattering-heavy directions, and promotes the pruning of suboptimal 3D Gaussians.
Training Objective
We adopt the photometric loss as follows:
$L_{\text{photo}}=(1-\lambda_{\text{s}})L_{\text{R-L1}}(I_{c},I_{gt})+\lambda_{\text{s}}L_{\text{R-SSIM}}(I_{c},I_{gt}),$ (22)
where $L_{\text{R-SSIM}}$ and $\lambda_{\text{s}}$ denote the Regularized SSIM loss (li2024watersplatting) and its corresponding weight, respectively, with $\lambda_{\text{s}}=0.2$. In addition to the baseline, we incorporate our proposed components into the final training objective:
$L_{\text{total}}=L_{\text{photo}}+\lambda_{\text{tri}}L_{\text{tri}}+\lambda_{\text{epi}}L_{\text{epi}}+\lambda_{\text{res}}L_{\text{res}},$ (23)
where $\lambda_{\text{tri}}$, $\lambda_{\text{epi}}$, and $\lambda_{\text{res}}$ are the weights of the proposed loss terms. We empirically set $\lambda_{\text{tri}}=0.1$ and $\lambda_{\text{res}}=0.01$, while $\lambda_{\text{epi}}$ is annealed from $0.4$ to $0.2$.
Experiment
Datasets and Metrics
We evaluate underwater scene reconstruction on four scenes from SeaThru-NeRF dataset (levy2023seathru) and two scenes from In-the-Wild dataset (tang2024neural). In addition, to evaluate scene restoration performance, we generate simulated underwater and foggy scenes using the Fern scene from the LLFF dataset (mildenhall2019local). We simulate underwater scenes following previous methods (levy2023seathru) using normalized depth obtained from a depth foundation model (yang2024depth) with $\beta_{D}=[1.3,\ 1.2,\ 0.9]$, $\beta_{B}=[0.95,\ 0.85,\ 0.7]$, and $\beta_{\infty}=[0.07,\ 0.2,\ 0.39]$. For foggy scenes, we use $\beta_{D}=[0.5,\ 0.5,\ 0.5]$ and $\beta_{B}=\beta_{\infty}=1.2$. To validate our method, we compare scene quality using PSNR, SSIM (wang2004image), and LPIPS (zhang2018unreasonable).
Implementation Details
Our method builds on Nerfstudio (tancik2023nerfstudio), and employs a TCNN-based MLP (tiny-cuda-nn). We use 7K/3K training steps for densification/finetuning on SeaThru-NeRF (levy2023seathru) and simulated datasets, and 10K/5K steps on In-the-Wild (tang2024neural). At each iteration, $b_{v}$ is sampled from [0.4,0.4][-0.4,0.4], and $b_{h}$ is set to $1.5b_{v}$. We enable $L_{\text{tri}}$ and $L_{\text{epi}}$ during iterations 4K–8K and 6K–12K for 10K and 15K training steps, respectively. The $\tau_{\alpha}$ for synthetic epipolar depth is set to 0.8. The input resolution is increased from $\times 1/4$ to $\times 1/2$ at $1/5$ of total training steps, and from $\times 1/2$ to full resolution at $2/5$ to boost the training process. For depth-aware alpha adjustment, $w$ is set to 0.5 for $t<t_{\alpha}$ and is decayed to 0 at $t_{\alpha}$. We set the transition step $t_{\alpha}$ to 4K and 10K for 10K and 15K training steps, respectively. All experiments use a single RTX A6000.
UW Rendered UW Restored Fog Rendered Fog Restored
Ours
Refer to caption Refer to caption Refer to caption Refer to caption
WaterSplatting
Refer to caption Refer to caption Refer to caption Refer to caption
SeaSplat
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of scene reconstruction and restoration in underwater and foggy environments.
Performance Comparison
Real-world Underwater Scenes
We evaluate OceanSplat with previous methods, including the official and Nerfstudio versions of SeaThru-NeRF (levy2023seathru), as well as 3DGS (kerbl20233d), SeaSplat (yang2024seasplat), and WaterSplatting (li2024watersplatting). All results were reproduced using the publicly released code. As shown in Table 1, OceanSplat achieves superior performance in novel view synthesis across most real-world underwater scenes. On the SeaThru-NeRF dataset (levy2023seathru), OceanSplat outperforms the previous state-of-the-art, WaterSplatting (li2024watersplatting), and SeaThru-NeRF-NS (levy2023seathru) by average PSNR margins of 1.05 dB and 2.88 dB, respectively. On the In-the-Wild dataset (tang2024neural), OceanSplat reduces the average LPIPS by $0.021$ compared to 3DGS and surpasses SeaSplat (yang2024seasplat) with average gains of $0.46$ dB in PSNR and $0.014$ in SSIM, along with a $0.024$ reduction in LPIPS.
Figure 4 compares RGB and depth renderings across three underwater scenes. In the Curaçao scene, prior methods suffer from degraded reconstructions due to floating artifacts, whereas our method models the 3D scene with significantly fewer floating artifacts and accurate object representation. In J.G-Redsea scene, prior methods yield plausible RGB images but their depth maps reveal erroneous 3D Gaussians in water regions. In contrast, our method maintains consistent geometry throughout. In the Coral scene, most prior methods (levy2023seathru; kerbl20233d; yang2024seasplat) fail to capture fine-grained coral structures, and WaterSplatting (li2024watersplatting) shows geometry loss and depth holes, while ours reconstructs a faithful 3D scene.
We additionally compare training time, inference FPS, and VRAM usage with prior methods. As shown in Table 3, our method trains faster and consumes less memory than SeaThru-NeRF (levy2023seathru), while remaining competitive among 3DGS-based approaches. Compared to WaterSplatting (li2024watersplatting), our method requires more rasterizations per iteration and least-squares solving for the synthetic epipolar depth prior, resulting in longer training time per-iteration. Nevertheless, it trains over $\times 2$ faster than SeaSplat (yang2024seasplat) while retaining comparable inference speed to other 3DGS-based methods.
Method Training $\downarrow$ FPS $\uparrow$ Memory (GB) $\downarrow$
SeaThru-NeRF 18 h 25 m 0.10 37.7
SeaThru-NeRF-NS 3 h 21 m 0.29 15.1
3DGS 22 m 150.28 3.7
SeaSplat 51 m 90.05 6.0
WaterSplatting 10 m 88.19 5.8
Ours 19 m 85.67 7.6
Table 3: Comparison of training time, rendering speed (FPS), and memory consumption with prior methods.
Simulated Scattering Scenes
We compare the scene reconstruction and restoration performance of SeaThru-NeRF-NS (levy2023seathru), SeaSplat (yang2024seasplat), WaterSplatting (li2024watersplatting), and our proposed OceanSplat using a simulated dataset. As shown in Table 2, for the novel view synthesis task, OceanSplat achieves an average PSNR improvement of 7.39 dB over SeaSplat (yang2024seasplat) and 0.67 dB over WaterSplatting (li2024watersplatting). For scene restoration, our method also demonstrates superior performance, yielding average SSIM gains of 0.048 over SeaSplat (yang2024seasplat) and 0.021 over WaterSplatting (li2024watersplatting). Figure 5 shows that, in both simulated scenes, previous methods assign 3D Gaussians to the scattering medium (e.g., ceilings or walls), resulting in restoration holes, whereas our method accurately represents geometric structures in both underwater and foggy scenes.
Configuration PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
Table 4: Ablation study demonstrating the effectiveness of each component in our method.
Configuration PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Training $\downarrow$
[…]
Table 5: Ablation study on stereo configurations, comparing the impact of horizontally and vertically translated camera poses on performance and training time.
Ablation Study
In Table 4, we evaluate the effectiveness of each component on the Curaçao scene. Removing $L_{\text{tri}}$ yields the largest PSNR drop, confirming that trinocular view consistency guides spatial optimization of 3D Gaussians, enabling structurally coherent novel view synthesis. Excluding $L_{\text{epi}}$ significantly increases LPIPS, validating our self-supervised depth regularization in underwater scenes with a lack of geometric cues. Omitting $L_{\text{res}}$ and depth-aware opacity $\alpha^{d}$ degrades overall quality, highlighting their role in mitigating medium entanglement through reduced depth ambiguity.
Table 5 compares rendering performance and training time across different stereo configurations, including the baseline (li2024watersplatting). Binocular stereo setup $(P_{c}+P_{h})$ improves over baseline, while adding another horizontal stereo view $(P_{c}+P_{h}+P_{h})$ provides minimal gain despite longer training. In contrast, introducing vertical disparity $(P_{c}+P_{h}+P_{v})$ achieves a $0.29$ dB increase in PSNR over the binocular setup. This shows that using horizontally and vertically translated camera poses enhances performance without compromising training time, as shown in Figure 3.
Conclusion
We introduce OceanSplat, a novel framework for enhancing geometric consistency in underwater scene reconstruction. Primarily, we propose trinocular view consistency across translated camera views along two orthogonal axes, effectively aligning 3D Gaussians with scene geometry. In addition, we introduce a synthetic epipolar depth prior derived via triangulation from translated camera views, enabling self-supervised geometric regularization. Lastly, we design a depth-aware alpha adjustment that uses depth and viewing direction to suppress medium-induced floating artifacts. Through these contributions, OceanSplat achieves high-fidelity underwater scene reconstruction. Future work will focus on addressing non-rigid object representations and eliminating the dependency on SfM, thereby broadening the applicability to dynamic underwater scenes.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2023-00217689, 50%; RS-2024-00358935, 40%), and in part by Institute of Information & communications Technology Planning & Evaluation (IITP) under the Artificial Intelligence Convergence Innovation Human Resources Development grant funded by the Korea government (MSIT) (IITP-2026-RS-2023-00254177, 10%). Minseong Kweon acknowledges travel support from the Minnesota Robotics Institute (MnRI).